OpenStudy (anonymous):

how does one simplify (solve for x) \[0\ge\frac{x^{2}-2x+3-4}{x^{2}-2x-3}\]

OpenStudy (anonymous):

ok so first you need to combine your like numbers so 3-4 and also x^2-2x

OpenStudy (anonymous):

I know 3-4 = -1, what's this about x^2-2x?

OpenStudy (anonymous):

basically what doodler said but even i dont understand that so as i was saying

OpenStudy (anonymous):

I got this:\[0\ge\frac{x^{2}-2x-1}{x^{2}-2x-3}\]

OpenStudy (anonymous):

And then this:\[0\ge\frac{x^{2}-2x-1}{(x-3)(x-1)}\]

OpenStudy (anonymous):

we cant square a letter so if we ignore the "^2" what do we have

OpenStudy (anonymous):

Top is unfactorable as far as I know

OpenStudy (anonymous):

try this lol

OpenStudy (anonymous):

"We can't square a letter" ...what?

OpenStudy (anonymous):

you have X^2 or commonly know as "X squared"

OpenStudy (anonymous):

\[\text{Aren't we solving for }x\text{ here?}\]

OpenStudy (anonymous):

That doesn't give me x^2-2x-1...?

OpenStudy (anonymous):

yes but we cant solve for X untill we can combine all of our like numbers

OpenStudy (anonymous):

They want to simplify it

OpenStudy (anonymous):

I did combine all the like values @IMANINJA127

OpenStudy (anonymous):

not your X values

OpenStudy (anonymous):

Well I meant solve for x

OpenStudy (anonymous):

And either way I'm doing the same thing aren't I?

OpenStudy (anonymous):

well you need to combine all of your X values first before you can solve

OpenStudy (anonymous):

I did...

OpenStudy (anonymous):

I got \[(x-2)(x-1)= x^{2}-2x-x+2=x^{2}-3x+2\]

OpenStudy (anonymous):

...I got lost

OpenStudy (anonymous):

alyways if u comebine all of your like numbers divide by the number of x etc 2x, 3x, 4x untill it is 1X= (unkown)

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

Combined form:\[0\ge\frac{x^{2}-2x-1}{x^{2}-2x-3}\]

OpenStudy (anonymous):

...Maybe I should just quadratic function this

OpenStudy (anonymous):

@Abhisar

OpenStudy (anonymous):

o-o

OpenStudy (anonymous):

calling in the big guns here

OpenStudy (anonymous):

I think I'll stick to @MDoodler 's explanation

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So is it \[(x-2)(x+1)?\]

OpenStudy (anonymous):

nevermind

OpenStudy (anonymous):

this might help http://www.webmath.com/redfract.html

OpenStudy (anonymous):

how do I put this huge equation into that tiny thing though

OpenStudy (anonymous):

@Demonx341 it's not fitting in

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

idk xD maby put it into parts.

OpenStudy (anonymous):

\[\text{*does not compute*}\]

OpenStudy (anonymous):

@MDoodler I got the answer you showed me using the quadratic formula, thanks :)